OpenStudy (rockycatsrule):

Solve the equation for y. 4x – 5y = 9 A.y= -4x-5 B. y=4/5 x -9/5 <---- this is the one i think is right but i'm not sure??? C. y= -5/4 x -5/9 <----- or maybe this one but still not sure yet???? D. y= 5/4 x + 9/4

TheSmartOne (thesmartone):

Hi

TheSmartOne (thesmartone):

4x – 5y = 9 Lets add 5y to both sides What do you get? :)

OpenStudy (rockycatsrule):

hey

OpenStudy (rushwr):

@rockycatsrule the answer is B that's right!

OpenStudy (rockycatsrule):

thanks

OpenStudy (rushwr):

4x-5y =9 4x-9 = 5y \[y = \frac{ 4x-9 }{ 5 } = \frac{ 4 }{ 5 }x - \frac{ 9 }{ 5 }\]

OpenStudy (rockycatsrule):

oh now i get what u did so thx so much.. so do u think u could help with some more

OpenStudy (rushwr):

yeah sure, no problem at all. But close this question and open a new one. :)

OpenStudy (rockycatsrule):

kk

OpenStudy (rushwr):

And also click on the blue colour best response button to medal the best answer :)